OpenStudy (anonymous):

Find the general solution of each differential function: dy/dx = 2xe^y

OpenStudy (ribhu):

same question runger

OpenStudy (anonymous):

no it isn't. Same instructions but I need help finding antiderivative of 1/e^y

OpenStudy (loser66):

-e^-y

OpenStudy (anonymous):

how did you get that?

OpenStudy (loser66):

take (-e^-y)'= e^(-y)

OpenStudy (loser66):

or just let -y =u and do something on e^u

OpenStudy (ribhu):

e^-y dy = 2x dx

OpenStudy (ribhu):

@runger723 it became like one of those which we did earlier

OpenStudy (loser66):

When I took cal1, I stuck at this problem since I tried to figure out the method to find out int of some tricky thing. My classmate gave me this method: try derivative of some simple thing to get the tricky one.

OpenStudy (loser66):

Like this, we know (e^y )' = e^y, and int (e^y) = e^y also how about e^(-y)? play around with -e^(-y) :)

OpenStudy (loser66):